OpenStudy (anonymous):

a 2000 kg truck is brakes on a stop in 15m from an initial speed of 12 m/s.how much force was required?

OpenStudy (ghazi):

v^2=U^2+2*a*S...so here final velocity =0 therefore substituing the value will give deceleration = 4.8m/s^2 and force = mass * acceleration therefore =2000*4.8=9600N

OpenStudy (anonymous):

thank you for the fast response sir Ghazi :)

OpenStudy (ghazi):

anytime